OpenStudy (anonymous):

find coordinates of x/ y intercept and max or mini point on the grapg of these quad fuction. show answer by obtaining reasoing with sumbolic forms help: 9a ) f(x)=x(7-x) d) f(x)=-(x-3)(x+5) g) f(x)=2(x-3)(x-8)

OpenStudy (phi):

what's with reposting this?

OpenStudy (anonymous):

my internet messed up

OpenStudy (phi):

Try this one f(x)=-(x-3)(x+5) replace x with a 0 f(0)= -(0-3)(0+5) can you do the arithmetic?

OpenStudy (anonymous):

i dont remember how to

OpenStudy (phi):

do the stuff in parens first (0-3) = ?

OpenStudy (anonymous):

-3?

OpenStudy (phi):

and (0+5) = ?

OpenStudy (anonymous):

5

OpenStudy (phi):

now -(-3)(5) multiply it out the - sign out front is the same as -1 -1* -3 * 5

OpenStudy (anonymous):

-15

OpenStudy (phi):

a minus times a minus is a plus

OpenStudy (anonymous):

so it would 15

OpenStudy (phi):

so you found f(0) and that is the y-intercept. now find the x - intercept. that is where f(x)= 0 or, in this case, -(x-3)(x+5) = 0 what x value makes this zero?

OpenStudy (anonymous):

f=15 for the y itnecept?

OpenStudy (phi):

y= 15 or f(0)= 15 is how you say it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so now how do i find x intecept

OpenStudy (phi):

now find the x - intercept. that is where f(x)= 0 or, in this case, -(x-3)(x+5) = 0 what x value makes this zero?

OpenStudy (anonymous):

3nd -5

OpenStudy (phi):

so those are your x-intercepts

OpenStudy (anonymous):

okay how would i write that

OpenStudy (phi):

The parabola peaks at exactly between the two x-intercepts so find the average of -5 and 3 to get where it peaks.

OpenStudy (anonymous):

how do i do that

OpenStudy (phi):

you could mark them down and start counting from -5 to 3 and then go to the half way point.

OpenStudy (phi):

or you could add them together and divide by 2

OpenStudy (anonymous):

i got -1

OpenStudy (phi):

next, we need to know is this a max or a min? Do you know? One way is put -1 in for x and figure out f(-1)

OpenStudy (anonymous):

can u write it out

OpenStudy (phi):

f(x)=-(x-3)(x+5) replace the x with a -1

OpenStudy (anonymous):

ok how to solve

OpenStudy (phi):

Here is a picture.